OpenStudy (anonymous):

Find the exact circumference, in terms of , of a circle with the given radius. 3¼ inches

OpenStudy (anonymous):

2* pi * radius

OpenStudy (anonymous):

It would be 20.4 correct?

OpenStudy (phi):

To be exact, you must say something like \( 2*3.25*\pi= 6.5 \pi \) As you know, pi is a number with unending decimals: 3.14159.... with no pattern if you round pi to 3.14 you will not get an "exact" answer, just a close one.